# Rigorous Function Calculi in Ariadne

### v2

@@ 5.3 Polynomial models
In Ariadne, unit polynomial models $p\pm e$ are defined by TaylorModel<ValidatedTag,F,FE>, where F is the type used for the coefficients of the polynomial and FE the type used for the error bound, The current implementation uses a sparse representation of the polynomial, with separate lists for the multi-indices $\alpha\in\mathbb{N}^{n}$ and the coefficients $c\in\mathbb{F}$. These are ordered by reverse lexicographic order, which makes inserting a constant term fast, and allows efficient evaluation using Horner’s rule. Alternative representations include a dense format, with only coefficients being stored, which may be useful for low-dimensional polynomials, and a representation using sparse storage for the multi-indices themselves, which may be useful for very high-dimensional polynomials.
@@ 5.4 Affine models
6 Solving Algebraic and Differential Equations
@@
6.1 Algebraic equations
Solvers for a simple algebraic equation $f(y)=0$ or a parameterised algebraic equation $f(x,y)=0$ with solution $y=h(x)$ are simply called Solver. Ariadne implements an IntervalNewtonSolver based on the interval Newton operator (Moore, 1966), and a KrawcykzSolver based on the related Krawcykz operator, which is more reliable but slower.
@@ 6.1 Algebraic equations
attempts to find a solution of the equation $f(x)=0$ in the box $D$, throwing a SolverException if no solution is found. The method delegates to an abstract method
SolverInterface::step(ValidatedVectorMultivariateFunction f,
@@
similar to (Berz and Hoefkens, 2001). If $f:\mathbb{R}^{n}\times\mathbb{R}\rightarrow\mathbb{R}$, then
@@ 6.3 Hybrid systems
We therefore see that the function calculus we have described provide exactly the building blocks needed to rigorously compute the evolution of a hybrid system. For more details, see (Collins et al., 2012) and the website (Ariadne, 2002-, ).
7 Examples
We now give examples of the use of Ariadne’s function calculus for solving algebraic and differential equations using the Python interface. We shall use as a running example the Fitzhugh-Nagumo system (Fitzhugh, 1961; Nagumo et al., 1962), which is defined by the equations
@@ 7.1 Fixed points
vw_rng=BoxDomainType([[dy_(1.0),dy_(1.5)],[dy_(0.75),dy_(1.25)]])
Finally, we use the implicit method to compute the parametrised fixed-point $(v(I_{\mathrm{ext}}),w(I_{\mathrm{ext}}))$:
@@
7.2 Differential equations
To solver the Fitzhugh-Nagumo system, we use an Integrator class. We first define the point and desired time interval:
@@ 7.2 Differential equations
flow_step = integrator.flow_step(f,vw0,h)
@@
flow_step = integrator.flow_step(f,vw0,h)
@@ 8.1 Alternative bases
8.2 $C^{1}$ function calculus
A $C^{1}$ function calculus includes uniform errors on both zeroth and first derivatives. We represent univariate $C^{1}$ function over a unit domain, $f:[-1\!:\!+1]\rightarrow\mathbb{R}$ by a polynomial $p$, and provide error bounds
@@ 8.3 Multivalued functions
8.4 Measurable functions
Since measurable functions are only defined up to equivalence almost everywhere, they cannot be computably evaluated. Instead, a general representation is given in terms of preimages of open sets, with $f^{-1}(V)=U_{\infty}$ where $U_{\infty}$ is a lower-measurable set (Collins, 2020a):
@@ 9 Conclusions
Work in the immediate future will focus on further improvements to the efficiency and accuracy of the tool. Partially implemented future extensions include function models in Chebyshev bases, a calculus for differentiable functions, and Weierstrass approximation of general continuous functions. Work is also in progress on more advanced classes of systems, notably on the evolution of nondeterministic hybrid systems described by differential inclusions (Z̆ivanović and Collins, 2010), and on model-checking linear temporal logic. Theoretical work is in progress on the evolution of stiff continuous dynamics, the analysis of stochastic systems.

### v4

@@ 5.3. Polynomial models
Unit polynomial models $p\pm e$ are defined by TaylorModel<ValidatedTag,F,FE>, where F is the type used for the coefficients of the polynomial and FE the type used for the error bound, The current implementation uses a sparse representation of the polynomial, with separate lists for the multi-indices $\alpha\in\mathbb{N}^{n}$ and the coefficients $c\in\mathbb{F}$. These are ordered by reverse lexicographic order, which makes inserting a constant term fast, and allows efficient evaluation using Horner’s rule. Alternative representations include a dense format, with only coefficients being stored, which may be useful for low-dimensional polynomials, and a representation using sparse storage for the multi-indices themselves, which may be useful for very high-dimensional polynomials.
@@ 5.4. Affine models
6. Solving Algebraic and Differential Equations
@@
6.1. Algebraic equations
Solvers for a simple algebraic equation $f(y)=0$ or a parameterised algebraic equation $f(x,y)=0$ with solution $y=h(x)$ are simply called Solver. Ariadne implements an IntervalNewtonSolver based on the interval Newton operator (Moo, 66), and a KrawcykzSolver based on the related Krawcykz operator Kra (69), which is more reliable but slower.
@@ 6.1. Algebraic equations
attempts to find a solution of the equation $f(x)=0$ in box $D$, throwing a SolverException if no solution is found. The method delegates to an abstract method
SolverInterface::
step(ValidatedVectorMultivariateFunction f,
@@
similar to (BH, 01). If $f:\mathbb{R}^{n}\times\mathbb{R}\rightarrow\mathbb{R}$, then
@@ 6.3. Hybrid systems
We therefore see that the function calculus we have described provide exactly the building blocks needed to rigorously compute the evolution of a hybrid system. For more details, see (CBGV, 12) and the website (Ariadne , ).
7. Examples
We now give examples of the use of Ariadne’s function calculus for solving algebraic and differential equations using the Python interface. We shall use as a running example the Fitzhugh-Nagumo system (Fit, 61; NAY, 62), which is defined by the equations
@@ 7.1. Fixed points
vw_rng=BoxDomainType([[dy_(1.0),dy_(1.5)],
[dy_(0.75),dy_(1.25)]])
Finally, we use the implicit method to compute the parametrised fixed-point $(v,w)$ as a function of $I_{\mathrm{ext}}$:
@@
7.2. Differential equations
To solve the Fitzhugh-Nagumo system, we use an Integrator class. We first define the point and a desired time interval:
@@ 7.2. Differential equations
flow_step = integrator.flow_step(f,vw0,suggest(h))
@@
flow_step = integrator.flow_step(f,vw0,suggest(h))
@@ 8.1. Alternative bases
8.2. Differentiable functions
A differentiable ($C^{1}$) function calculus includes uniform errors on both zeroth and first derivatives. We represent univariate $C^{1}$ function over a unit domain, $f:[-1\!:\!+1]\rightarrow\mathbb{R}$ by a polynomial $p$, and provide error bounds
@@ 8.3. Multivalued functions
8.4. Measurable functions
Since measurable functions are only defined up to equivalence almost everywhere, they cannot be computably evaluated. Instead, a general representation is given in terms of preimages of open sets, with $f^{-1}(V)=U_{\infty}$ where $U_{\infty}$ is a lower-measurable set (Col20b, ):
@@ 9. Conclusions
Work in the immediate future will focus on further improvements to the efficiency and accuracy of the tool. Partially implemented future extensions include function models in Chebyshev bases, a calculus for differentiable functions, and Weierstrass approximation of general continuous functions. Work is also in progress on more advanced classes of systems, notably on the evolution of nondeterministic hybrid systems described by differential inclusions (Z̆C, 10), and on model-checking linear temporal logic. Theoretical work is in progress on the evolution of stiff continuous dynamics, the analysis of stochastic systems.